What is the frequency of a photon with a wavelength of 781 nm. Report your answers to three significant digits.
1
Expert's answer
2020-11-12T07:12:33-0500

Solution:

the frequency ƒ = the speed of light c / the wavelenght λ

ƒ = 3.0×108 m/s × 7.81×10-7 m = 3.84 ×1014 Hz  

Answer: ƒ = 3.84 ×1014 Hz